OpenStudy (rizags):

0=(3)(a^2)(b)(a+b) where a= 2^x - 4 and b= 4^x - 2 What is the sum of all real numbers x that satisfy

OpenStudy (kagıtucak):

3 possiblities a^2=0 , b=0 , or a+b=0

OpenStudy (kagıtucak):

if a^2=0 a=0 so 2^x=4 x=2

OpenStudy (kagıtucak):

if b=0 4^x=2 x=1/2

OpenStudy (kagıtucak):

ifa+b=0 4^x+2^x-6=0 let 2^x=c then c^2+c-6=0 (c+3)(c-2)=0 c=-3 c=2 2^x=-3 this is impossible 2^x=2 x=1

OpenStudy (kagıtucak):

all x are 2,1/2,1 sum=7/2

OpenStudy (rizags):

so 7/2?

OpenStudy (rizags):

THAnk you SOo much I understand it

OpenStudy (kagıtucak):

yes 7/2 x=2,1,1/2 nop